40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 2 21:22:07 2015, rana, koji, Configuration, IOO, AOM inserted in beam and aligned 
    Reply  Fri Sep 4 10:25:19 2015, Steve, Configuration, IOO, AOM stage is ready stage_AOM-EOM.jpgA-EOMholder.pdf
       Reply  Mon Sep 7 18:25:16 2015, rana, Configuration, IOO, AOM stage is ready 
Message ID: 11343     Entry time: Tue Jun 2 21:22:07 2015     Reply to this: 11576
Author: rana, koji 
Type: Configuration 
Category: IOO 
Subject: AOM inserted in beam and aligned 

We spent an hour today to put the AOM back in the beam before the PMC and verified that the diffraction is working.

  1. The fuse holder was missing from the rack. We inserted a 5A fuse. We expect that the quiesscent draw is < 0.5 A. The power is from the +24V Sorensen supply.
  2. The alignment was tricky, but we optimized it as well as we could in translation and the RZ direction. Its a fixed mount still.
  3. We noticed that according to the datasheet, the polarization is wrong! It wants S-Pol light and we're giving it P-Pol. How come no one noticed this? We expect that the efficiency is reduced because of this. We (Steve) need to brainstorm what kind of mount we can use there to mount it at 90 deg to the plane of the table.
  4. The lens after the AOM has f = +400 mm. The distance from the AOM to the lens is ~800-900 mm so its not so terrible. However, if someone were to put the AOM halfway between the turning mirrors there, the beam diffraction would be canceled.
  5. The AOM input impedance seems to be 50 Ohm as advertised. The previous Koji entry claim of 25 Ohm is mysterious. We checked the Ohmage by sending a signal into the AM input of the AOM using the DS345 which as a 50 Ohm output. 1 Vpp from the DS345 made 1 Vpp on the input of the AM input as measured by Oscope connected by T with high impedance setting.
  6. With 0.5 V offset and a 1 Vpp signal, we get ~20-25% modulation of the power.sad
  7. We have left it running with a 4444.4 Hz modulation and a small amplitude. This is to see if we can use this to measure the cavity poles of the MC and the arms.
  8. We noticed some hash on the Teed input monitor. It was backstreaming of the RF drive. Whoever uses this thing in an ISS feedback ought to make sure to put an RF choke between the servo and this AOM driver.

We also removed a 50/50 pickoff mirror which was used to take one of the NPRO -> EOM polarizer reject beams and send it across the table into a floppy dump. Its now hitting a closer floppy dump. Let's stop using these crappy anodized aluminum flappers anywhere, Steve.

We also noticed that the PMC REFL path uses a W1 from CVI to send the PMC reflection to the REFL RFPD. The dim beam from the AR coated surface is being used rather than the bright beam from the uncoated surface. Ooops. Steve, can you please order another W1 for 1064 from CVI, but get it with a 2-3 deg wedge angle? This one has a wedge which is too small.

ELOG V3.1.3-